HUBBLE SPACE TELESCOPE DAILY REPORT #3101 PERIOD COVERED: 0000Z (UTC) 04/23/02 - 0000Z (UTC) 04/24/02 Daily Status Report as of 114/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed STIS/CCD 9139 (Variability in the UV Spectrum of 3C 279: Testing Models for the Gamma-Ray Emission in Blazars) The Space Telescope Imaging Spectrograph (CCD) was used to determine whether gamma- rays are produced on the scale of the broad line region, using multi-epoch UV snapshots of 3C 279 to find and characterize its LyAlpha variability. This will validate the mirror model for generating the intense gamma-ray emission from blazars. 3C 279 is the best-studied blazar from radio to gamma-ray wavelengths, showing frequent large flares over days to months. Variability in the UV continuum and LyAlpha will directly reveal any coupling between the jet ionizing flux and the broad-line region emission, providing clues to the physics and energetics of all radio-loud AGN. All observations completed without reported incident. 1.2 Completed Three Sets of ACS/WFC 9575 (Default {Archival} Pure Parallel Program) The Advanced Camera for Surveys (WFC) was used to test ACS pure parallels in POMS. There were no reported problems. 1.3 Completed Seven Sets of WF/PC-2 9318 (POMS Test Proposal: WFII Parallel Archive Proposal Continuation) The WF/PC-2 was used to perform the generic target version of the WFPC2 Archival Pure Parallel program. The program was used to take parallel images of random areas of the sky, following the recommendations of the Parallels Working Group. There were no reported problems. 1.4 Completed Eleven Sets of NICMOS/1/2/3 8945 (Cooling System Monitoring) The Near Infrared Camera and Multi-Object Camera (NIC1, NIC2 and NIC3) was used 1) to measure NICMOS detector performance during the cool-down and steady state operation of the NCS. 2}. to demonstrate stability {+/-0.1K} of the NICMOS detector temperature at the optimal science operating temperature. and 3}.to demonstrate repeatability {+/-0.1 K} of NICMOS detector temperature following changes from the optimal science operating temperature. No problems were encountered. 1.5 Completed Three Sets of WF/PC-2 8937 (Cycle 9 Supplemental Darks pt2/3) The WF/PC-2 was used obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. No problems were encountered. 1.6 Completed Two Sets of STIS/CCD 8904 (Bias Monitor-Part 2) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the bias in the 1x1, 1x2, 2x1, and 2x2 bin settings at gain=1, and 1x1 at gain = 4 in order to build up high-S/N superbiases and track the evolution of hot columns. The proposal completed with no anomalous activity. 1.7 Completed Five Sets of ACS/WFC/HRC 8947 (Weekly Test) The Advanced Camera for Surveys (WFC and HRC) was used to perform basic tests to monitor, the read noise, the development of hot pixels and test for any source of noise in ACS CCD detectors. This program will be executed at least once a day for the entire lifetime of ACS. There were no problems reported. 1.8 Completed STIS/CCD 8902 (Dark Monitor-Part 2) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the darks. No problems were reported. 1.9 Completed Two Sets of STIS/CCD 9496 (UV Spectroscopic Investigation of Any Bright, Newly Discovered Comet) The Space Telescope Imaging Spectrograph (CCD) was used to perform a Target of Opportunity program to investigate any bright comet {mboxV<=sssim7} that is newly discovered during Cycle 11, including comets of any dynamical class. The main scientific objective is to obtain accurate abundance measurements for several key cometary species: CO from the CO 4PG bands, cotwo from the CO Cameron bands, stwo from the stwo B-X bands, cstwo from CS emissions, and water from OH emissions. The UV Cameron band emission is currently the only way to probe cotwo in comets. No anomalous activity occurred. 1.10 Completed STIS/CCD/MA1 9067 (UV Detectability of Bright Quasars in the Sloan Fields) The Space Telescope Imaging Spectrograph (CCD and MA1) was used to take MAMA spectra of approximately 30 new, bright, high-redshift quasars in each of the next three cycles. The observations completed with no reported problems. 1.11 Completed STIS/MA2 9573 (NUV-MAMA Daily Dark Monitor) The Space Telescope Imaging Spectrograph (MA2) was used to perform daily monitoring of the NUV MAMA detector dark noise in order to monitor the effects of thermal changes on the NUV dark rate. No problems occurred. 1.12 Completed STIS/CCD 9110 (A Search for Kuiper Belt Object Satellites) The Space Telescope Imaging Spectrograph (CCD) was used to investigate whether the large number of collisions thought to have taken place in the primordial Kuiper belt suggest that many Kuiper belt objects {KBOs} could have suffered binary-forming collisions similar to that which formed the Pluto -- Charon binary. Detection of such KBO satellites would allow measurement of KBO masses, would help to understand the past collisional environment of the Kuiper belt, and would give a context to the otherwise unique-seeming formation of the Pluto -- Charon binary. The proposal completed with no reported problems. 1.13 Completed Two Sets of STIS/MA2 8920 (Cycle 10 MAMA Dark Measurements) The Space Telescope Imaging Spectrograph (MA2) was used to perform the routine monitoring of the MAMA detector dark noise, and is the primary means of checking on health of the MAMA detectors systems through frequent monitoring of the background count rate. The proposal completed with no reported anomalous activity. 1.14 Completed STIS/CCD 9163 (Kinematics Of Emission-Line Gas Disks In Radio-Quiet Galaxies) The Space Telescope Imaging Spectrograph (CCD) was used to measure central black hole {BH} masses from the rotation rate of the emission-line gas, and to determine the nature and structure of the gas disks. Three galaxies with no radio jets and with Halpha+[NII] emission have been identified, which have dust disks similar to those commonly seen in our sample of radio-loud active galaxies. All observations completed nominally. 1.15 Completed FGS/1 9168 (The Distances to AM CVn Stars) Fine Guidance Sensor (FGS) #1R was used to determine the parallaxes and proper motions of the five brightest of the seven known AM CVn systems. AM CVn systems are binaries where mass is transferred from a completely hydrogen-deficient, degenerate mass donor to a white dwarf primary through a helium accretion disk. A better understanding of these systems is crucial for a number of reasons: (1) to study the late stages of binary evolution, (2) to study the effect of chemical composition on the physics of accretion discs, (3) to estimate their contribution to the Supernovae Ia rate, and (4) to estimate their contribution to the gravitational radiation background. All observations completed with no reported problems. 1.16 Completed Two Sets of STIS/CCD 9317 (Pure Parallel Imaging Program: Cycle 10) The Space Telescope Imaging Spectrograph (CCD) was used to perform the default archival pure parallel program for STIS during cycle 10. There were no reported problems. 1.17 Completed STIS/CCD 9139 (Variability in the UV Spectrum of 3C 279: Testing Models for the Gamma-Ray Emission in Blazars) The Space Telescope Imaging Spectrograph (CCD) was used to determine whether gamma- rays are produced on the scale of the broad line region, using multi-epoch UV snapshots of 3C 279 to find and characterize its LyAlpha variability. This will validate the mirror model for generating the intense gamma-ray emission from blazars. 3C 279 is the best-studied blazar from radio to gamma-ray wavelengths, showing frequent large flares over days to months. Variability in the UV continuum and LyAlpha will directly reveal any coupling between the jet ionizing flux and the broad-line region emission, providing clues to the physics and energetics of all radio-loud AGN. All observations completed without reported incident. 1.18 Completed Five Sets of NICMOS/1/2/3 8944 (Filter Wheel/Mechanisms Mini-Functional Test) The Near Infrared Camera and Multi-Object Camera (NIC1, NIC2 and NIC3) was used to perform an early engineering test to verify the aliveness, functionality, operability, and electro-mechanical calibration of the NICMOS filter wheel motors and assembly. This was the first use of the NICMOS filter wheel mechanisms since they were disabled by ground command in January, 1999. This test was designed to obviate concerns over possible deformation or breakage of the fitter wheel "soda-straw" shafts due to excess rotational drag torque and/or bending moments which may be imparted due to changes in the dewar metrology from warm-up/cool-down. No anomalous incident occurred. 1.19 Completed Three Sets of ACS/WFC/HRC 9075 (Cosmological Parameters from Type Ia Supernovae at High Redshift) The Advanced Camera for Surveys (WFC and HRC) was used to obtain a Hubble diagram of Type Ia supernovae {SNe Ia} that will be of long lasting value as a record of the expansion history of the universe. No problems were encountered. 1.20 Completed Five Sets of WF/PC-2 8941 (Cycle 10 UV Earthflats) The WF/PC-2 was used to monitor flat field stability by obtaining sequences of earth streak flats to improve the quality of pipeline flat fields for the WFPC2 UV filter set. The proposal had no problems. 1.21 Completed WF/PC-2 8940 (Cycle 10 Earth Flats) The WF/PC-2 was used to monitor flatfield stability by obtaining sequences of Earth streak flats to construct high quality flat fields for the WF/PC-2 filter set. These flat fields will allow mapping of the OTA illumination pattern and will be used in conjunction with previous internal and external flats to generate new pipeline superflats. The proposal completed without incident. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 9 Successful: 9 The acquisition at 114/084229Z defaulted to fine lock backup on FGS-2 only when the scan step limit for FGS-1 was exceeded. HSTAR 8626 was written. Three proposals occurring during the period of the next report, STIS 9051, NICMOS 8945, and WF/PC-2 8941, may have been affected. Scheduled Re-acquisitions: 4 Successful: 4 2.2 FHST Updates: Scheduled: 33 Successful: 33 2.3 Operations Notes: Operations requests were utilized to set up the tasking orders for the NICMOS filter wheel test at 113/1207Z and to establish the NICMOS memory monitors for the test at 113/1214Z. Then, at 113/1548Z, another operations request enabled the filter wheel move. Between 113/1717Z and 113/1738Z, the filter wheels were moved and data was collected. Initial analysis of the data indicated that the mechanism voltages and current during the moves was as expected and very close to the last on-orbit test in 1997. There was no significant temperature change. During the moves, the voltage dropped approximately 0.3V on all three wheels and the current spiked by 0.75A. There were no retries on filter wheel 1, and that wheel moved within 14 counts of its commanded position. The return to blank was within 16 counts of the commanded position. There was one retry on filter wheel 2. This is not anomalous or cause for concern. The position was off by 126 counts (the tolerance is 110 counts) which caused the retry. On the second attempt, the filter wheel reached its position within nine counts. There was an expected overshoot of four steps. There were no retries on filter wheel 3. That wheel moved 60 steps as expected. The return to blank was 64 steps. The final position was within 58 counts of the commanded position. This test will continue over the next several days. The NCC PID control was returned to the weighted neon temperatures from the NICMOS dewar temperature at 113/1755Z, per an operations request. Using ROP SR-9a, the SSR-3 EDAC error counter was cleared at 113/1929Z. The SSR-1 EDAC error counter was cleared at 113/2110Z per ROP SR-1A. The SSR-3 playback pointer was set at 114/0348Z as directed by ROP SR-3. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of Servicing Mission Orbital Verification and the gradual resumption of normal science observations and calibrations. /CAW